The July 3rd Clermont visit yielded interesting colony status. We were able to inspect all the colonies. All but two have reasonable brood conditions.
One of the two marginal colonies has an opened supersedure queen cell located on the face of a frame and several charged but not yet capped queen cells. The examining team reported seeing eggs. There were capped drone cells in a pattern that could indicate the queen had failed and the bees have replaced her. This would be a reasonable assumption and the desirable outcome.
The other marginal colony had no brood so it will be carefully watched and likely supplemented with a queen-right nuc at our next visit. This colony had reasonable bee numbers, just no brood.
One ten-frame colony was compacted to a three-super stack and the other ten-frame colony down to a four-super stack. Both were five supers before this change. The frames that were removed were older with some going back to before the hives were brought to Clermont. Both have been at Clermont for the five years the training yard has existed. Neither of these hives are overly strong and there is ample evidence of mite issues. There is brood in both, but pin holes in some caps and a spotty brood pattern were apparent. After compacting, there appeared to be appropriate bee cover.
The balance of the colonies are reasonable in strength with solid brood in a normal mid-summer arrangement. These colonies are in a mix of eight-frame and five-frame equipment. Most of the eight-frame colonies are four supers which is the preferred Clermont overwinter configuration.
The five-frame (nuc) colonies are a mix of two-super hives, a few three-super hives and several one-super hives. The targeted Clermont overwinter configuration will be three supers with a two-frame feeder providing thirteen frames for the bees to use.
The shorter eight-frame colonies will be brought up to four supers. All the colonies will have a two-frame internal feeder installed lowering the total frame count by two. Depending on late July and August conditions, supplemental syrup and protein feeding will be initiated. The pollen substitute will be fed externally.
The next visit (scheduled for July 17th) will be to treat all the colonies for mites. Because the mid-July daytime temperature is normally above safe formic acid limits and oxalic acid does not treat for mites in capped cells, Apivar will be used. An additional fall treatment will be considered.
Summer nuc production was discussed. Details are being worked out. It is possible that this could occur on July 31st. Additional information will be provided when it becomes available.